OpenStudy (pucusana):

Pythagorean Triples are integer solutions to the Pythagorean Theorem, a^2+b^2=c^2. As such, the solutions identified in the above four problems all qualify as Pythagorean Triples. Using what you know about similar triangles and the four Pythagorean triples you have identified, complete the following table: Leg1 Leg2 Hypotenuse 3 4 6 8 9 12 5 12 10 24 15 36 7 25 14 50 21 75 15 17 30 34 45 51

OpenStudy (pucusana):

can you help me?

OpenStudy (nethercreep333):

probs not sorrry @ParthKohli @Qwertty123 @IrishBoy123 they might can

OpenStudy (pucusana):

can you help me, please?

OpenStudy (qwertty123):

Hello! :D

OpenStudy (irishboy123):

just compute the sides, and scale up the integer multiples so a 3,4,5 justifies a 6, 8, 10

OpenStudy (pucusana):

what about leg 2?

OpenStudy (pucusana):

hello are you there?

OpenStudy (irishboy123):

which one is that?

OpenStudy (pucusana):

the second on the list, the middle

OpenStudy (mathstudent55):

@pucusana Use the Pythagorean theorem. If the legs are 3 and 4, what is the hypotenuse?

OpenStudy (irishboy123):

oh i see take \(a^2 + b^ 2 = c^2 \) and re-write as \(a^2 = c^2 - b^ 2\) is that the confusion?

OpenStudy (pucusana):

thank you for the tip you all have been very kind. :)